1) write the equation of the circle with center at(2,1)and which passes thru the point (2,7)
2) write the equation of the circle with center at (-1,-2) and which passes thru the point(-1,2)

Answer by venugopalramana(3286)

1) write the equation of the circle with center at(2,1)and which passes thru the point (2,7)
STD.EQN.OF CIRCLE WITH CENTRE AS (H,K) AND RADIUS = R IS
(X-H)^2 + (Y-K)^2 =R^2.............H=2.........K=1..........SO
EQN.IS
(X-2)^2+(Y-1)^2=R^2..IT PASSES THROUGH (2,7)..HENCE
(2-2)^2+(7-1)^2=R^2=36
HENCE EQN.OF CIRLE IS
(X-2)^2+(Y-1)^2=36

2) write the equation of the circle with center at (-1,-2) and which passes thru the point(-1,2)
DO BY SAME METHOD AS ABOVE